# The DECADE cosmic shear project III: validation of analysis pipeline using spatially inhomogeneous data

### v1

@@ Appendix B Addition results from Survey Property splits
First, Figure 9 shows the footprint of the area-based splits used in Figure 4. The different colors distinguish the area spanned by a given subset and its complement. The splits on airmass are highly correlated with those on DCR, while there are similar correlations between the splits on exposure time, exposure count, and magnitude limit, and also between the two Milky Way properties (star density and extinction). All maps are shown at $\texttt{NSIDE}=1024$, which is the resolution used to make the splits.
@@ Appendix C Validation of shape noise term in the covariance matrix model
Figure 11: Ratios of the shape noise component in the covariance matrix between that measured from 600 simulations and analytical predictions ($\sigma^{\rm ana}_{\rm SN}$). We show the square root of the diagonal of the covariance matrix for each bin pair, and for two scenarios: (1) where we use the galaxy positions that are identical to the data and contain some level of spatial inhomogeneity (blue, $\sigma^{\rm sim}_{\rm rand}$) and (2) where we reposition each galaxy randomly inside the footprint to create a homogeneously distributed sample (red, $\sigma^{\rm sim}_{\rm rand,uni}$). Both sets of points scatter within 5%, indicating our analytical prescription of shape noise is accurate. We further find that the two sets of points do not show any qualitatively different behaviors, indicating the the spatial inhomogeneity in the galaxy number counts does not affect the shape noise component of the covariance significantly. Uncertainties on the simulation covariance are obtained from a leave-one-out jackknife across the 600 simulations.
@@
As mentioned in Section 4.2, the $\mathbf{C}_{\rm SN}$ term includes the effect of the survey mask but not that of the variation in galaxy counts within said mask. However, we now show that the latter effect does not cause an appreciable shift in the covariance matrix estimates.
@@
This validation is done by randomly rotating the measured ellipticities of the DECADE shape catalog and computing the $\xi_{\pm}$ data vectors using TreeCorr (Jarvis et al. 2004). Such techniques have been used, for example, by Troxel et al. (2018) for validating the shape noise model. By repeating this rotation step for $\mathcal{O}(10^{3})$ different realizations, we obtain a data-driven covariance matrix of $\xi_{\pm}$ that includes only shape noise, but accounts for the relevant spatial variations in the data by directly using the positions and shapes from the DECADE catalog. We also perform a variant where the data vector is computed after removing the inhomogeneities, simply by reassigning DECADE galaxies to new positions that now uniformly sample the survey mask. Figure 11 shows the ratio between the shape noise terms from the simulation-based method relative to those from the analytic method. This comparison is done for both the fiducial simulation-based covariance as well as the one with randomized galaxy positions; note that the analytic estimate is the same in either comparison. We find that in both cases, the analytic method matches the simulated one to within $5\%$. The fact that both sets of points are consistently within $5\%$ of the analytic estimate highlights that the spatial variation of the galaxy number density is a negligible factor in calculating the $\mathbf{C}_{\rm SN}$ term. The errorbars on the simulation covariance matrices are estimated using a leave-one-out jackknife, i.e., we remove one simulated data vector from the full set and generate the covariance matrix.
@@
Finally, we have also compared the analytical covariance matrix (using all terms in Equation 7) with a jackknife-based covariance matrix computed on the data. The main additional aspect that this checks — beyond the test above — is any significant discrepancy in the off-diagonal terms of the covariance matrix. However, the relative differences of the two estimates depends on the cosmological parameters assumed for the analytic estimate and therefore cannot be a rigorous test before unblinding our cosmology constraints. In addition, the jackknife method is known to be biased on large scales where the variation across jackknife regions is not fully Gaussian; we have also verified this limitation via simulated data vectors. As such, we performed this comparison only as an approximate check and do not show the results here given the limitations in their interpretability.

### v2

@@ Appendix B Additional results from catalog splits
First, Figure 9 shows the survey footprint for each of the area-based splits that was analysed and presented earlier in Figure 4. The different colors distinguish the area spanned by a given subset and its complement. The splits on airmass are highly correlated with those on DCR, while there are similar correlations between the splits on exposure time, exposure count, and magnitude limit, and also between the two Milky Way properties (star density and extinction). All maps are shown at $\texttt{NSIDE}=1024$, which is the resolution used to make the splits.
@@
Next, Table 4 and Table 5 list the constraints on two IA parameters — $a_{2}$, the amplitude of the tidal torquing term, and $\eta_{2}$, the redshift dependence of this amplitude — as well as the $\chi^{2}$ of the maximum posterior point in our chains alongside its associated p-value. The Seeing+ result in Table 4 is our worst-fit datavector, with $p=0.002$. This barely passes our criteria of $p>0.0015(3\sigma)$ discussed in Appendix C of Paper IV. However, we remain unconcerned as all estimates of the p-value do not include any corrections for the “look elsewhere” effect. In brief, we have split the same data into forty-six different subsets. The distribution of p-values will no longer follows a standard null distribution. The exact deviations can be quantified using a simulation-based approach, where one generates $N$ simulated datasets, measures datavectors for the forty-six subsets on it, and computes the resulting p-values. This calibration is unfeasible in our case as the simulations must have a high enough fidelity to reproduce all properties of the data (e.g., the connection between airmass and object properties), and we must generate at least $N\sim 100$ realizations, which results in $4600$ tests/chains. Returning to the origin of this discussion — the low p-value of the datavector best-fit for Seeing+ — we note that Figure 13 in Paper I finds no correlation between the mean shear and seeing (which is referred to in that Figure 13 as “fwhm”). This indicates the seeing is not contaminating our measured shear.
Table 4 and Table 5 also lists the IA parameters of each result. This is motivated by our results in Paper IV, where we find a non-zero amplitude for the tidal torquing term of the TATT IA model. We find here that the results of a given split and its complement are in good consistency with one another. The largest deviation is between the DECADE East and West split, but is consistent within $2.4\sigma$. This is still completely consistent with just statistical fluctuations, and we reemphasize our point above about the “look elsewhere” effect, which can impact the parameter posteriors as well and generally boosts the measured significance. We also note that the $a_{2}$ constraints between the objects with $T/T_{\rm psf}\gtrless 1$ are consistent within $0.9\sigma$, which is a clean test of the impact of PSF systematics on this constraint as the amplitude of such systematics generally scales linearly with this ratio (Paulin-Henriksson et al. 2008). This measurement of minimal impact from PSF contamination is consistent with our results in Figure 7, which show the minimal impact on derived cosmology constraints.
Finally, we address the impact of correlations on the consistency between constraints of a subset and of its complement. As mentioned before, the shape noise in the measurement from a subset and from its complement are completely uncorrelated as no source galaxy is shared between the two. However, the two subsets’ galaxy samples can still cover similar (or the same) sky area and therefore have correlations due to cosmic variance. These correlations are generally subdominant relative to shape noise but we perform a quantitative test of this here. To do so, we generate 800 mock catalogs from the CosmoGrid simulation suite (Kacprzak et al. 2023), using the same procedure as Yamamoto & Becker et al. (2025, see their Section 2.2). We generate these catalogs for each subset, while accounting for the exact sky-area mask of the subset and for the $n(z)$ of all tomographic bins in that subset. This generates, for each simulation realization, a catalog for every subset of interest. The resulting set of catalogs can be used to study the correlation between the cosmology constraints of a subset and its complement.
We perform this test for two subsets: Maglim±, where the sky-area mask is quite patchy (and therefore the two subsets can be more strongly correlated by cosmic variance), and the $m_{r}\gtrless 22.5$ split, where the catalogs have significantly different $n(z)$ but share the same sky area. As a result, we generate four variant datavectors per simulation realization. For each datavector (so, for each subset in each of the 800 realizations), we estimate the cosmology constraints using the maximum likelihood optimizer available in Cosmosis. [^12] We use a maximum likelihood as our point estimate of $S_{8}$ instead of generating a full bayesian posterior and taking the mean/median as the latter is computationally infeasible for this task. The former is a common approach for quoting constraints in lensing datasets (e.g., Secco & Samuroff et al. 2022; Amon et al. 2022; DES and KiDS Collaborations et al. 2023). We then compute the correlation between the 800 estimates of $S_{8}$ from a subset and from its complement. For Maglim±, we see a correlation of $\rho=0.18$ between the estimates and for $m_{r}\gtrless 22.5$, we see a correlation of $\rho=0.28$. These are relatively mild correlations and only change the inferred parameter-shift significances in Section 6 by 10-15%. The correlations between a subset and its complement for other split definitions will be lower than those estimated above. The Maglim± is the patchiest split of all considered (see Figure 9) and the $m_{r}\gtrless 22.5$ selection is highly correlated with the other selections on signal-to-noise, brightness, size etc. used in Figure 6. In summary, the mild correlations between the constraints from a subset and its complement only cause minor ($\sim 10\%$) changes to the inferred significances of the parameter shifts between the two constraints, and do not change any of our qualitative discussion. Critically, all split tests continue to pass comfortably even if we account for this minor correlation.
@@ Appendix C Validation of shape noise term in the covariance matrix model
Figure 11: Ratios of the shape noise component in the covariance matrix between that measured from 600 simulations ($\sigma^{\rm sim}_{\rm SN}$) and analytical predictions ($\sigma^{\rm ana}_{\rm SN}$). We show the square root of the diagonal of the covariance matrix for each tomographic bin pair, and for two scenarios: (1) where we use the galaxy positions that are identical to the data and contain some level of spatial inhomogeneity (blue, $\sigma^{\rm sim}_{\rm rand}$) and (2) where we reposition each galaxy randomly inside the footprint to create a homogeneously distributed sample (red, $\sigma^{\rm sim}_{\rm rand,uni}$). Both sets of points scatter within 5%, indicating that our analytical prescription of shape noise is accurate. We further find that the two sets of points do not show any qualitatively different behaviors, indicating that the spatial inhomogeneity in the galaxy number counts does not affect the shape noise component of the covariance significantly. Uncertainties on the simulation covariance are obtained from a leave-one-out jackknife across the 600 simulations.
@@
As mentioned in Section 4.2, the $\mathbf{C}_{\rm SN}$ term of the covariance matrix $\mathbf{C}$ includes the effect of the survey mask but not that of the variation in galaxy counts within said mask. However, we now show that the latter effect does not cause an appreciable shift in the covariance matrix estimates.
@@
This validation is done by randomly rotating the measured ellipticities of the DECADE shape catalog and computing the $\xi_{\pm}$ data vectors using TreeCorr (Jarvis et al. 2004). Such techniques have been used by Troxel et al. (2018) for validating the shape noise model. By repeating this rotation step for $\mathcal{O}(10^{3})$ different realizations, we obtain a data-driven covariance matrix of $\xi_{\pm}$ that includes only shape noise, but accounts for the relevant spatial variations in the data by directly using the positions and shapes from the DECADE catalog. We also perform a variant where the data vector is computed after removing the inhomogeneities, simply by reassigning DECADE galaxies to new positions that now uniformly sample the survey mask. Figure 11 shows the ratio between the shape noise terms from the simulation-based method relative to those from the analytic method. This comparison is done for both the fiducial simulation-based covariance and the one with randomized galaxy positions; note that the analytic estimate is the same in either comparison. We find that in both cases, the analytic method matches the simulated one to within $5\%$. The fact that both sets of points are consistently within $5\%$ of the analytic estimate highlights that the spatial variation of the galaxy number density is a negligible factor in calculating the $\mathbf{C}_{\rm SN}$ term. The errorbars on the simulation covariance matrices are estimated using a leave-one-out jackknife method, i.e., we remove one simulated data vector from the full set and generate the covariance matrix.
@@
Finally, we have also compared the analytical covariance matrix (using all terms in Equation 9) with a jackknife-based covariance matrix computed on the data. The main additional aspect that this checks — beyond the test above — is any significant discrepancy in the off-diagonal terms of the covariance matrix. However, the relative differences of the two estimates depend on the cosmological parameters assumed for the analytic estimate and therefore cannot be a rigorous test before unblinding our cosmology constraints. In addition, the jackknife method is known to be biased on large scales where the variation across jackknife regions is no longer independent (e.g., Shirasaki et al. 2017); we have also verified this limitation via simulated data vectors. As such, we performed this comparison only as an approximate check and do not show the results here given the limitations in their interpretability.